1). Suppose the demand and supply for gas generators in a coastal Florida town are as follows:

Qd = 1800 – 3P

Qs = 2P – 200.

a. Find the equilibrium price and quantity of generators.

b. Find the consumer surplus, producer surplus, and total gains from trade at the equilibrium price.

c. Question 2 asks you to consider the economic impacts of “price gouging.” Price gouging refers to suppliers raising their prices on essential goods (bottled water, gas, generators, airline tickets) during a crisis. Using the links that follow or a Google search, describe the costs and benefits of price gouging as described by economists. Then, discuss your opinion on the issue.
Answer #1

a) Set QD=QS

1800-3P = 2P-200

1800+200 = 2P+3P

2000 = 5P

P = 2000/5 = 400

Q = 1800-3*400 = 600

b) Maximum reservation price = 1800/3 = 600

Minimum reservation price = 100

CS = 0.5*600*(600-400) = 60000

PS = 0.5*600*(400-100) = 90000

TS = CS+PS = 60000+90000 = 150000